# Mining Legal Arguments to Study Judicial Formalism

### v1

@@ 4.4 MADON dataset for NLP experiments
For the experiments described in the upcoming section, we split the MADON dataset into train (70%), validation (20%), and test (10%) sets. The split was stratified to proportionally represent holistic labels and issuing court throughout the subsets, as summarized in Table 4.
@@
5 Experiments
5.1 Experiment setup and tasks
We design three different experiments to investigate the connection between court document texts, arguments, and formalism in the MADON dataset. In general, mining arguments is often split into several subtasks [eger-etal-2017-neural] and we follow this pattern by defining our subtasks into the following: As an initial probe whether models are capable of detecting arguments at all, we train it to predict the presence and absence of arguments in a paragraph. We do not distinguish between a single argument and multiple arguments and focus only on detecting the presence. This experiment design also serves as a sanity check to see if our models are able to distinguish argumentative and non-argumentative paragraphs. The second and most difficult task requires the extraction of the exact argument types from a given paragraph. Since there can be more than one argument in a paragraph, we formulate the problem as a multilabel classification problem with a binary representation for the presence of each argument type separately. If there is no argument present, the output should be ‘False’ for each argument type. Finally, we tackle the question of formalism for the entire document. This third task should also provide us with insights whether models are capable of understanding the document holistically without having additional labels. Here we experiment with different approaches and pipelines; the simplest approach passes the entire text of the decision to the model and performs a binary classification (formalistic/non-formalistic). In the following, we refer to our experiments as:
Task 1: Argument presence detection
We classify paragraphs into argumentative and non-argumentative based on the presence or absence of any argument in it.
Task 2: Argument type classification
Entire paragraphs are classified in a multilabel setting into the eight types of traditional legal argument types described in Section 3.2.
Task 3: Holistic formalism classification
We classify entire judicial decisions as formalistic or non-formalistic.

### v2

@@ 5.2 Data for NLP experiments
For the experiments, we split the MADON dataset into train (70%), validation (20%), and test (10%) sets. The split was stratified to proportionally represent holistic labels and issuing court throughout the subsets, as summarized in Table 3.
In addition, we compiled a new corpus for continued pretraining of our models. We collected and preprocessed 300,511 decisions from the Czech Supreme Court (152,925) and all administrative courts (147,586), comprising the Supreme Administrative Court (74,937) and eight Regional Administrative Courts (72,631). We make the corpus available to the AI&Law research community. [^6] https://github.com/trusthlt/madon The decisions were collected from the official anonymized databases of the respective courts. Our corpus complements existing Czech legal datasets, namely [novotna2019czechcourtdecisionscorpus] and [Paulik_2025]. Compared to novotna2019czechcourtdecisionscorpus, we provide regional administrative court decisions, richer metadata, scraping scripts, and, most importantly, coverage of the six year period after their cut-off date, i.e., 09/2018-09/2024. The dataset from Paulik_2025 focuses on the Constitutional Court, a different institution not covered here.